# On A Scale From 1 to 5: Quantifying Hallucination in Faithfulness Evaluation

### v2

@@ 4.1 Scoring Quality
Figures 2(a) - 2(e) elaborates on the residuals of the scores. Overall, GPT-4 with zero-shot rubrics performed the best among all models, and Llama3.1-70B followed closely after. The tuned HHEM-extrinsic model also shows high accuracy with a slight flaw in scoring sentence-level JsonTG dataset. A slight performance drop can be observed transitioning from the gold hypothesis (2(a)) to its sentence level (2(b)). GPT-4 scored 6.8% out of 1393 gold sentences to "1-highly unfaithful", which was mainly due to incomplete information in individual sentences. As a result, several unfaithful occurrences were assigned reason 1 in Table 3. Furthermore, JsonTG reference seems to be consistently challenging to all models across all tasks. After inspecting the LLM provided reasoning, we found that this is because our adjective verification rubrics to capture "adjectives without any basis" contradicts the JsonTG generation task which allows LLMs to come up with expressive adjectives. For posterity, it is best to adapt the rubrics to a per-use case basis for better accuracy.
LLMs with fewer parameters such as Llama3.1-8B and Claude3-Haiku seem to struggle to understand the rubrics. Llama3.1-8B tends to be overly strict in that it scored all tasks to "3-slightly unfaithful" and below, indicating the hypothesis is unfaithful. This seems to be a result of Llama3.1-8B’s following the instructions too literally. In addition, Llama3.1-8B seems to have confused "factual consistency" with "granularity". The example LLM reasoning can be found on row 2 and 3 in Table 3. In contrast, Haiku tends to be extremely lenient. It failed consistently on both hallucination tasks while scoring almost perfectly on the gold hypothesis. It appears that Haiku was able to capture the key information in the hypothesis, but could not make the association to the contradictory part in the reference. Row 4 in Table 3 illustrated an example where Haiku incorrectly scored an intrinsic hypothesis to "5-highly faithful".
It is interesting that the above observation is not reflected in the overall residual plot in Figure 2(e) as the scores got smoothed out by half of the tasks with good performance. However, it should be noted that although the models’ strictness and leniency are consistent across different tasks, in reality, it would be difficult to integrate them into the production pipeline due to the unpredictability brought by their scoring bias.
@@
Figure 3: Scores on a hypothesis with different hallucination percentage. (top) Intrinsic hallucination. (bottom) Extrinsic hallucination. The scores range from "1-highly unfaithful" to "5-highly faithful". The percentage is simulated by gradually increasing the amount of hallucinating sentences in the 5-sentence gold hypothesis (51 samples). Models with an "SO" suffix indicate "score only", in other words, the LLMs were not asked to provide reasoning. Overall, models perform better towards intrinsic hallucination. Including reasoning in response can increase the sensitivity of the model in picking up hallucinations.
@@
Unlike LLMs, HHEM models do not experience significant performance degradation from full length to sentence-level hypothesis. This is likely because the base Deberta model was originally trained for NLI, which uses a slightly fuzzy definition for "textual entailment" that a premise entails a hypothesis given a very probable inference can be made from the premise (Dagan et al., 2005). Thus, a piece of a sentence from the hypothesis would receive a higher score during the prediction as long as the encoded information can be inferred from the source. It should be noted that HHEM models are naturally prone to input token size, evaluating on sentence level could improve efficiency by inducing fewer segments on the reference.
@@
HHEM models are also less sensitive toward extrinsic hallucinations as compared to intrinsic hallucinations. Fine-tuning on extrinsic synthetic JSON data alone helped HHEM-extrinsic model gain better performance on almost all tasks. However, during fine-tuning, we noticed that it was rather easy to over-tune the HHEM model and made it less tolerant to any type of creative contents. We first tried tuning with 400 each gold and hallucination examples for 5 epochs, which resulted in overly strict scores. By reducing the number of epochs to 3, we were able to achieve the performance as shown in Figure 2. We believe if the model is trained on more diverse synthetic examples with a carefully crafted dataset, the results would be more stable than what we have shown here.
@@
4.2 Scoring Progression
Figures 3 illustrates the score progression trends with respect to the percentage of the hallucinating information. To recap, the percentage-based hallucination dataset was created by randomly replacing the N-numbers of gold sentences with hallucinating ones. Each hallucinating sentence contains a subtle change as demonstrated in Table 2. Consequently, the hallucination percentage is in fact the percentage of sentences containing unfaithful information. In this iteration, we also tested LLM’s scoring capability when the "reasoning" component was removed from its response. The score-only model’s performance is suffixed with "SO".
As shown in Figures 3(a) - 3(b) for both intrinsic and extrinsic hallucination, the faithfulness score decreases accordingly as the percentage of hallucination increases from 0% to 100%. This trend is consistent even for lower-performing LLMs such as Llama3-8B and Claude3-Haiku. As mentioned earlier, all models tend to perform better on intrinsic hallucination over extrinsic ones. We can potentially explain this observation further by comparing the changes in score going from lower to higher hallucination percentage. Taking GPT-4 as an example, in Figure 3(a), when the intrinsic hallucination percentage increases from 0% -> 20% -> 40%, the score decreases from 4.90 -> 3.23 -> 2.26; whereas for extrinsic hallucination, the score decreases in a more gradual manner, namely 4.90 -> 4.13 -> 3.15, with every step taken. GPT-4’s sensitivity to identify the initial 20% of intrinsic hallucination is roughly mapped to detecting 40% of extrinsic hallucination. Tracing back to our rubrics, although "avoiding speculative or creative content generation" was included in the guidelines, the score assignment was lenient towards extrinsic hallucination. We designed the "3-slightly unfaithful" to allow unverifiable information to be present without negatively impacting the reader’s experience. Tightening the rubrics could help increase the sensitivity in detecting extrinsic hallucinations. However, one should proceed with caution and support the rubric’s development by using similar experiments to ensure that the LLMs maintain a good tolerance between unfaithful and creative content.
We also quantified the scores without reasoning by simply removing the instruction to justify the reason for grading. Overall, in Figures 3(a) - 3(b), all LLMs demonstrated various levels of performance degradation. For both intrinsic and extrinsic hallucinations, scoring without reasoning made the LLMs more lenient towards the unfaithful content. As the percentage of hallucination increases, the decreases in score converge much slower. As a result, none of the LLMs scored a 100% hallucinating content to "1-highly unfaithful". Another interesting observation is that the strictness of Llama3.1-8B and the leniency of Claude3-Haiku have both relaxed slightly without reasoning, which caused a performance "improvement". We argue that this is because the reasoning is the key to keeping the model "sharp", the "improvement" is really just a form of unresponsiveness when the "thinking" step is removed from the process.
@@ 4.3 Latency & Usage
Table 4 illustrates the latency and price for each model. All LLMs were tested through HTTP endpoints hosted through the same proxy service. The average time in seconds, number of output tokens, and prices were calculated in a sequential API call setting. By nature, smaller encoder models are inherently faster than decoder-based LLMs, thus we only computed a batch inference latency for HHEM models ran on an AWS G5 instance. The pricing referred to OpenAI  [^2] https://openai.com/api/pricing/ and AWS Bedrock  [^3] https://aws.amazon.com/bedrock/pricing/ standards. While GPT-4 yields the best accuracy, it is also the most costly model on both latency and price. On the other hand, although Llama3.1-70B seems to have balanced out the accuracy vs. latency, Llama3.1-70B provided reasoning can be barely useful to a human auditor due to little granularity. It should be noted that one could customize the prompt instruction to adjust the level of granularity to achieve the desired balance. For real-world usage, we recommend combining a fast model with acceptable performance as a first pass evaluation; then apply more powerful LLMs with detailed explanation to assist human audits.
@@ 5 Conclusion
We proposed a rubrics-based template to use LLMs to score faithfulness in NLG. We evaluated 5 popular LLMs on 4 travel-domain indutry dataset. The results indicate that in a zero-shot setting LLMs, e.g. GPT-4, outperforms NLI models, e.g. HHEM, on capturing both intrinsic and extrinsic hallucinations. HHEM model with fine-tuning on synthetic data can outperform LLMs in domain specific evaluation given carefully crafted training data. Lastly, we introduced a heuristic to quantify faithfulness score by hallucination percentage, which helps visualize the sensitivity when developing faithfulness evaluation models. Our future work include exploring the score association between different LLMs from the same family and tuning the rubrics to adjust the scoring sensitivity.
@@ 6 Limitations
    Due to resource limitations, we focused on testing only travel-domain dataset, but did not include popular open source faithfulness dataset. We justify that real-world dataset are often more challenging than research domain dataset, as there exist more noise in real-world data. We believe our approach can be generalized to other domains.
@@
    Due to privacy concerns, we plan to release a limited portion of the dataset. We’re in the process of legal review to determine which sections of the dataset can be made public.
@@
    Due to resource limitations, we were not able to evaluate a larger model from the Anthropics family, such as the Claude3.5-Sonnet (Anthropic, 2024b). It was not our intention to make an unfair comparison using only a smaller Claude model.
7 Acknowledgement
This work was conducted at and funded by Data and AI, Expedia Group Inc. The authors would like to thank the anonymous reviewers for their detailed, kind, and constructive reviews; Anirudh Kamalapuram Muralidhar for the initial work on the LLM rubrics; Magdalena Labori and Ananth Raj GV on the valuable feedback; Saman Enayati, Zoe Yang, and Mani Najmabadi for their supports; and David Cohen, Rose James, and Christina Riggs on legal guidance.
@@ A.1 Prompt Templates
All prompts in this paper was experimented with temperature set to 0 for consistency.
Table 5 contains the prompt used for LLM faithfulness grading. For score only grading, we simply removed the "reasoning" field from the output format. During our experiments, without enforcing controlled decoding to ensure consistent JSON output, we noticed that only GPT-4 and GPT-4o were able to always respond with a stringified JSON. Both Llama3.1-8B and Claude-Haiku tends to add heading and tailing texts and sometimes include unhashed double quotes which resulted in a JSON decoding challenge. For open source models which can be served by VLLM (Kwon et al., 2023), we recommend using VLLM’s controlled decoding strategy to formulate the structured output.
Table 6 illustrates prompts used for generating the synthetic intrinsic and extrinsic hallucination data. Each prompt follows the "Template" format as noted on row 1. Depending on the use case and the nature of the dataset, we customize the case-by-case instruction for the best results as denoted on row 2-7. It should be noted that a general prompt for generating synthetic data could work, however, the data generated might not be tricky enough so that the changes are subtle. For example, when our source data contains the following ground-truth: "3-bedroom" and "2-bathroom". There is a chance that a LLM confuses the numeric values and writes "2-bedroom and 2-bathroom" in the generation. By including explicit details in the instructions, we were able to mimic similar cases.
@@ A.2 Additional Analysis on Synthetic Data
Table 7 demonstrates several tricky intrinsic examples in the synthetic dataset. Prior studies have shown that language models tend to struggle with following the instructions when negations present in prompts or generations (Truong et al., 2023; Varshney et al., 2024). While a large portion of negations similar to row 1 were scored correctly, we did see that GPT-4 consistently fail on cases like row 2. In this case, the gold summary was in fact extracted word by word from the source conversation: "If your original card charge is still processing, it will be dropped automatically." The word "not" was obviously ignored every time when evaluating the intrinsic sentence against the source. Row 3 and 4 are good examples of near realistic synthetic data. Especially row 4 when the information is almost correct but not 100% accurate.
Table 8 showcases a few extrinsic hallucination examples generated by GPT-4o. We found some of the synthetic sentences extremely humorous, especially when the knowledge is added in an analogy form which potentially triggers a punchline. While these might be too "easy" to capture when served as extrinsic hallucinations, we thought it is worth pointing out this discovery, as it may be useful to humor research with LLMs.

### v3

@@ 4.1 Scoring Quality
Figures 2(a) - 2(e) elaborate on the residuals of the scores. Overall, GPT-4 with zero-shot rubrics performed the best among all models, and Llama3.1-70B followed closely thereafter. The tuned HHEM-extrinsic model also shows high accuracy with a slight flaw in scoring the sentence-level JsonTG dataset. A slight performance drop can be observed transitioning from the gold hypothesis (2(a)) to its sentence level (2(b)). GPT-4 scored 6.8% out of 1393 gold sentences to "1-highly unfaithful", which was mainly due to incomplete information in individual sentences. As a result, several unfaithful occurrences were assigned reason 1 in Table 3. Furthermore, JsonTG reference seems to be consistently challenging for all models across all tasks. After inspecting the LLM provided reasoning, we found that this is because our adjective verification rubrics to capture "adjectives without any basis" contradict the JsonTG generation task, which allows LLMs to come up with expressive adjectives. For posterity, it is best to adapt the rubrics to a per-use case basis for better accuracy.
LLMs with fewer parameters, such as Llama3.1-8B and Claude3-Haiku, seem to struggle to understand the rubrics. Llama3.1-8B tends to be overly strict in that it scored all tasks at "3-slightly unfaithful" and below, indicating the hypothesis is unfaithful. This seems to be a result of Llama3.1-8B’s following the instructions too literally. In addition, Llama3.1-8B seems to have confused "factual consistency" with "granularity". The example LLM reasoning can be found in rows 2 and 3 in Table 3. In contrast, Haiku tends to be extremely lenient. It failed consistently on both hallucination tasks while scoring almost perfectly on the gold hypothesis. It appears that Haiku was able to capture the key information in the hypothesis but could not make the association to the contradictory part in the reference. Row 4 in Table 3 illustrated an example in which Haiku incorrectly scored an intrinsic hypothesis as "5-highly faithful".
It is interesting that the above observation is not reflected in the overall residual plot in Figure 2(e), as the scores were smoothed out by half of the tasks with good performance. However, it should be noted that although the models’ strictness and leniency are consistent across different tasks, in reality, it would be difficult to integrate them into the production pipeline due to the unpredictability brought about by their scoring bias.
@@
Figure 3: Score progression on different hallucination percentage. (top) Intrinsic hallucination. (bottom) Extrinsic hallucination. The scores range from "1-highly unfaithful" to "5-highly faithful". The subset contains 51 samples with only 5-sentence hypothesis. Starting with the gold hypothesis (0%), the hallucination percentage is increased by 20% at every step, until all sentences are hallucinating sentences (100%). The "SO" suffix indicates "score-only" results, which the prompt was adjusted to exclude the "reasoning" in the output. When the hypothesis contains 0% hallucination, the expected score should be "5-highly faithful"; and when the hypothesis contains 100% hallucination, the expected score should be "1-highly unfaithful".
@@
Unlike LLMs, HHEM models do not experience significant performance degradation from full-length to sentence-level hypothesis. This is likely because the base Deberta model was originally trained for NLI, which uses a slightly fuzzy definition for "textual entailment" that a premise entails a hypothesis given a very probable inference can be made from the premise (Dagan et al., 2005). Thus, a piece of a sentence from the hypothesis would receive a higher score during the prediction as long as the encoded information can be inferred from the source. It should be noted that HHEM models are naturally prone to input token size, and evaluating on sentence level could improve efficiency by inducing fewer segments on the reference.
@@
HHEM models are also less sensitive toward extrinsic hallucinations as compared to intrinsic hallucinations. Fine-tuning on extrinsic synthetic JSON data alone helped the HHEM-extrinsic model to achieve better performance on almost all tasks. However, during fine-tuning, we noticed that it was rather easy to over-tune the HHEM model and made it less tolerant to any type of creative content. We first tried tuning with 400 each gold and hallucination examples for 5 epochs, which resulted in overly strict scores. By reducing the number of epochs to 3, we were able to achieve the performance as shown in Figure 2. We believe that if the model is trained on more diverse synthetic examples with a carefully crafted dataset, the results would be more stable than what we have shown here.
4.2 Scoring Progression
Figures 3 illustrate the progression trends of the scores with respect to the percentage of hallucination. Table 4 further shows the sensitivity of each model towards changes in amount of hallucination. To recap, the percentage-based hallucination dataset was created by randomly replacing the N-numbers of gold sentences with hallucinating ones. Each hallucinating sentence contains a subtle change, as demonstrated in Table 2. Consequently, the hallucination percentage is, in fact, the percentage of sentences containing unfaithful information. A gold sample is considered to contain 0% hallucination while a sample with only hallucinating sentences is considered to have 100% hallucination. In this iteration, we also tested the LLM scoring capability when the "reasoning" component was removed from its response. The score-only model’s performance is suffixed with "SO".
hallucination 0% 100% delta per 0% 100% delta per
percentage intrinsic intrinsic step (20%) extrinsic extrinsic step (20%)
GPT-4 4.90 1.06 -0.77 4.90 1.29 -0.72
Llama3.1-70B 4.92 1.12 -0.76 4.92 1.27 -0.73
Llama3.1-70B SO 4.62 1.12 -0.70 4.62 1.22 -0.68
GPT-4o-mini SO 4.78 1.39 -0.68 4.78 2.55 -0.45
GPT-4o-mini 4.35 1.00 -0.67 4.35 1.55 -0.56
GPT-4 SO 4.84 1.57 -0.65 4.84 2.16 -0.54
HHEM-extrinsic 4.16 1.15 -0.60 4.16 1.08 -0.61
HHEM-intrinsic 3.83 1.10 -0.55 3.83 1.96 -0.38
HHEM 4.25 1.55 -0.54 4.25 3.27 -0.33
Llama3.1-8B SO 3.06 1.08 -0.40 3.06 1.20 -0.37
Llama3.1-8B 2.62 1.07 -0.31 2.61 1.96 -0.28
Claude3-Haiku SO 4.61 3.67 -0.19 4.61 3.82 -0.16
Claude3-Haiku 5.00 4.65 -0.07 5.00 4.35 -0.13
Table 4: Model sensitivity vs. hallucination percentage. The larger the change (delta per step) is, the more sensitive the model is towards the amount of hallucination. In other words, GPT-4 is the best at detecting degrees of intrinsic hallucination, whereas Llama3.1-70B is the best at differencing levels of extrinsic hallucination.
As shown in Figures 3(a) - 3(b) for both intrinsic and extrinsic hallucinations, the faithfulness score decreases accordingly as the percentage of hallucination increases from 0% to 100%. This trend is consistent even for lower-performing LLMs such as Llama3-8B and Claude3-Haiku. As mentioned earlier, all models tend to perform better in intrinsic hallucinations than in extrinsic ones. We can potentially explain this observation further by comparing the changes in the score from a lower to a higher percentage of hallucinations. Taking GPT-4 as an example, in Figure 3(a), when the intrinsic hallucination percentage increases from 0% -> 20% -> 40%, the score decreases from 4.90 -> 3.23 -> 2.26; whereas for extrinsic hallucination, the score decreases in a more gradual manner, namely 4.90 -> 4.13 -> 3.15, with every step taken. GPT-4’s sensitivity to identify the initial 20% of intrinsic hallucination is roughly mapped to detecting 40% of extrinsic hallucination. Tracing back to our rubrics, although "avoiding speculative or creative content generation" was included in the guidelines, the score assignment was lenient toward extrinsic hallucination. We designed the "3-slightly unfaithful" to allow unverifiable information to be present without negatively impacting the reader’s experience. Tightening the rubrics could help increase the sensitivity in detecting extrinsic hallucinations. However, one should proceed with caution and support the rubric’s development by using similar experiments to ensure that the LLMs maintain a good tolerance between unfaithful and creative content.
We also quantified the scores without reasoning by simply removing the instruction to justify the reason for the grading. Overall, in Figures 3(a) - 3(b), all LLMs demonstrated various levels of performance degradation. For both intrinsic and extrinsic hallucinations, scoring without reasoning made the LLM more lenient toward unfaithful content. As the percentage of hallucinations increases, the decreases in the score converge much slower. The overall decrease in the "delta per step" in Table 4 further justifies that models are less sensitive to extrinsic hallucinations compared to intrinsic ones. Another interesting observation is that the strictness of Llama3.1-8B and the leniency of Claude3-Haiku have both relaxed slightly without reasoning, which caused a performance "improvement". We argue that this is because the reasoning is the key to keeping the model "sharp", the "improvement" is really just a form of unresponsiveness when the "thinking" step is removed from the process.
@@ 4.3 Latency & Usage
Table 5 illustrates the latency and price of each model. All LLMs were tested through HTTP endpoints hosted through the same proxy service. The average time in seconds, the number of output tokens and the prices were calculated in a sequential API call setting. By nature, smaller encoder models are inherently faster than decoder-based LLMs; thus we only computed a batch-inference latency for HHEM models run on an AWS G5 instance. The pricing referred to OpenAI  [^2] https://openai.com/api/pricing/ and AWS Bedrock  [^3] https://aws.amazon.com/bedrock/pricing/ standards. While GPT-4 yields the best accuracy, it is also the most costly model in both latency and price. On the other hand, although Llama3.1-70B seems to have balanced out the accuracy vs. latency, Llama3.1-70B provided reasoning can be barely useful to a human auditor due to little granularity. It should be noted that one could customize the prompt instruction to adjust the level of granularity to achieve the desired balance. For real-world usage, we recommend combining a fast model with acceptable performance as a first-pass evaluation; then apply more powerful LLMs with detailed explanation to assist human audits.
@@ 5 Conclusion
We proposed a rubric-based template to use LLMs to score faithfulness in NLG. We evaluated 5 popular LLMs on 4 travel-domain indutry dataset. The results indicate that in a zero-shot setting LLMs, e.g. GPT-4, outperform NLI models, e.g. HHEM, on capturing both intrinsic and extrinsic hallucinations. HHEM model with fine-tuning on synthetic data can outperform LLMs in domain-specific evaluation given carefully crafted training data. Lastly, we introduced a heuristic to quantify the faithfulness score over degrees of hallucination percentage, which helps to visualize the model’s sensitivity. Our future work includes exploring the score association between different LLMs of the same family and tuning the rubrics to adjust the scoring sensitivity.
@@ 6 Limitations
    Due to resource limitations, we focused on testing only travel-domain dataset, but did not include popular open source faithfulness dataset. We justify that real-world datasets are often more challenging than research domain datasets, as there exists more noise in real-world data. We believe that our approach can be generalized to other domains.
@@
    Due to resource limitations, we were unable to evaluate a larger model from the Anthropics family, such as the Claude3.5-Sonnet (Anthropic, 2024b). It was not our intention to make an unfair comparison using only a smaller Claude model.
@@
    Due to privacy concerns, we are in the process of legal review to determine which sections of the dataset can be made public.
7 Acknowledgment
This work was conducted by Data and AI, Expedia Group Inc. The authors would like to thank the anonymous reviewers for their detailed, kind and constructive reviews; Anirudh Kamalapuram Muralidhar for the initial work, Adam Johns and the team for the related work on the LLM rubrics; Magdalena Labori and Ananth Raj GV for the valuable feedback; Saman Enayati, Zoe Yang, and Mani Najmabadi for their supports; and David Cohen, Rose James, and Christina Riggs for legal guidance.
@@ A.1 Prompt Templates
All the prompts in this paper were experimented with a temperature set to 0 for consistency.
Table 6 contains the prompt used for LLM faithfulness grading. For score-only grading, we simply removed the "reasoning" field from the output format. During our experiments, without enforcing controlled decoding to ensure consistent JSON output, we noticed that only GPT-4 and GPT-4o were able to always respond with a stringified JSON. Both Llama3.1-8B and Claude-Haiku tend to add heading and tailing texts and sometimes include unhashed double quotes, which resulted in a JSON decoding challenge. For open source models which can be served by VLLM (Kwon et al., 2023), we recommend using VLLM’s controlled decoding strategy to formulate the structured output.
Table 7 illustrates prompts used for generating the synthetic intrinsic and extrinsic hallucination data. Each prompt follows the "Template" format as noted in row 1. Depending on the use case and the nature of the dataset, we customize the case-by-case instruction for the best results, as indicated in rows 2-7. It should be noted that a general prompt for generating synthetic data could work; however, the data generated might not be tricky enough, so that the changes are subtle. For example, when our source data contain the following ground truth: "3-bedroom" and "2-bathroom". There is a chance that an LLM confuses the numeric values and writes "2 bedrooms and 2 bathrooms" in the generation. By including explicit details in the instructions, we were able to mimic similar cases.
@@ A.2 Additional Analysis on Synthetic Data
Table 8 demonstrates several tricky intrinsic examples in the synthetic dataset. Previous studies have shown that language models tend to struggle to follow instructions when there are negations present in the prompts or generations (Truong et al., 2023; Varshney et al., 2024). Although a large portion of negations similar to row 1 were scored correctly, we did see that GPT-4 consistently fail on cases like row 2. In this case, the gold summary was, in fact, extracted word by word from the source conversation: "If your original card charge is still processing, it will be dropped automatically." The word "not" was obviously ignored every time when evaluating the intrinsic sentence against the source. Rows 3 and 4 are good examples of near-realistic synthetic data. Especially in row 4 where the information is almost correct, but not 100% accurate.
Table 9 presents a few examples of extrinsic hallucinations generated by GPT-4o. We found some of the synthetic sentences extremely humorous, especially when the knowledge is added in an analogy form which potentially triggers a punchline. While these might be too "easy" to capture when served as extrinsic hallucinations, we thought it is worth pointing out this discovery, as it may be useful to humor research with LLMs.